Data was collected for 300 fish from the North Atlantic. The length of the fish (in mm) is summarized in the GFDT below.

Lengths (mm)      Frequency   
180 - 189 1
190 - 199 16
200 - 209 71
210 - 219 108
220 - 229 83
230 - 239 18
240 - 249 3

What is the lower class boundary for the first class?

class boundary =
Answer #1

The lower class boundary for the first class is calculated as follows:=

The formula for the lower class boundary is ,

Lower Class Boundary=Lower Class Limit - width/2

Here Lower class limit for first class = 180

Width = upper limit - lower limit =189 - 180 = 9

Hence ,  Lower Class Boundary=Lower Class Limit - width/2

= 180 - 9/2

=180 - 4.5

Lower Class Boundary = 175.5